Find out which of the following functions from R to R are (i) One-to-one, (ii) Onto, (iii) One-to-one corre￾spondence.

(a)

f: R—>R defined by f(x) = x

(b)

f: R—>R defined by f(x) = |x|

(c)

f: R—>R defined by f(x) = x + 1

(d)

f: R—>R defined by f(x) = x^2

(e)

f: R—>R defined by f(x) = x^3

(f)

f: R—>R defined by f(x) = x – x^2

(g)

f: R—>R defined by f(x) = Floor(x)

(h)

f: R—>R defined by f(x) = Ceiling(x)

(i)

f: R—>R defined by f(x) = – 3x+4

(j)

f: R—>R defined by f(x)= – 3x^2 +7


1
Expert's answer
2021-07-01T12:15:48-0400

(a)

If "y_1=y_2," then "x_1=x_2=>" One-to-one.

For every "y\in \R" exists "x\in \R" such that "x=y=>" Onto

"f(x)=x" is One-to-one correspondence (a bijection).


(b)

"|-2|=2=|2|"

If "y=-2," then there is no "x\in \R" such that "|x|=-2."

"f(x)=|x|" is neither One-to-one nor Onto.


(c)

If "y_1=y_2," then "x_1+1=x_2+1=>x_1=x_2=>" One-to-one.

For every "y\in \R" exists "x\in \R" such that "x=y-1=>" Onto

"f(x)=x+1" is One-to-one correspondence (a bijection).


(d)


"(-2)^2=4=(2)^2"


If "y=-4," then there is no "x\in \R" such that "x^2=-4."

"f(x)=x^2" is neither One-to-one nor Onto.


(e)

If "y_1=y_2," then "x_1^3=x_2^3=>(x_1-x_2)(x_1^2+x_1x_2+x_2^2)=0"

"=>x_1=x_2=>" One-to-one.

For every "y\in \R" exists "x\in \R" such that "x=\sqrt[3]{y}=>" Onto

"f(x)=x^3" is One-to-one correspondence (a bijection).


(f)


"0-(0)^2=0=1-(1)^2"

If "y=1," then there is no "x\in \R" such that "x-x^2=1."

"f(x)=x-x^2" is neither One-to-one nor Onto.


(g)


"Floor(0.2)=0=Floor(0.5)"


If "y=0.5," then there is no "x\in \R" such that "Floor(x)=0.5."

"f(x)=Floor(x)" is neither One-to-one nor Onto.


(h)


"Ceiling(0.2)=1=Ceiling(0.5)"


If "y=0.5," then there is no "x\in \R" such that "Ceiling(x)=0.5."

"f(x)=Ceiling(x)" is neither One-to-one nor Onto.


(i)

If "y_1=y_2," then "-3x_1+4=-3x_2+4=>-3x_1=-3x_2"

"=>x_1=x_2=>" One-to-one.

For every "y\in \R" exists "x\in \R" such that "x=\dfrac{-y+4}{3}=>" Onto

"f(x)=-3x+4" is One-to-one correspondence (a bijection).


(j)


"-3(-1)^2+7=4=-3(1)^2+7"


If "y=8," then there is no "x\in \R" such that "x^2=8."

"f(x)=-3x^2+7" is neither One-to-one nor Onto.